A Complete Bibliography of Publications in the Proceedings of the American Philosophical

Society (1800–1849)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 August 2019 Version 1.00

Title word cross-reference

1 [Ano39j, Ano39p, Ano39t, Ano40u, Ano41t, Ano41-30, Ano42e, Ano42j, Ano43y, Ano44p, Ano44t, Ano46q, Ano47l, Ano47v, Ano48h]. 10 [Ano43l]. 12 [Ano43p]. 13 [Ano43e, Ano48d]. 14 [Ano45e]. 15 [Ano38q, Ano39k, Ano39q, Ano39u, Ano40v, Ano41u, Ano41-31, Ano42k, Ano42i, Ano42s, Ano43x, Ano43-40, Ano44q, Ano44u, Ano45i, Ano46r, Ano47m, Ano47w, Ano48i, Ano48x, Ano49k]. 16 [Ano38l, Ano38r, Ano38v, Ano39g, Ano40y, Ano41m, Ano41v, Ano42m, Ano42-29, Ano43-31, Ano44j, Ano45d, Ano45t, Ano46k, Ano46u, Ano47e, Ano47n, Ano48o, Ano49f, Ano49l]. 16th [Ano44g]. 17 [Ano38i, Ano39n, Ano39r, Ano40h, Ano40o, Ano40q, Ano41q, Ano41-32, Ano42t, Ano43z, Ano43-32, Ano43-36, Ano44r, Ano45n, Ano45x, Ano46d, Ano46m, Ano47h, Ano47x, Ano48p, Ano48t, Ano49e]. 18 [Ano38t, Ano39l, Ano39v, Ano40k, Ano40-27, Ano41w, Ano42o, Ano42u, Ano42z, Ano43w, Ano44v, Ano45g, Ano45p, Ano46g, Ano46w, Ano47o, Ano48g, Ano48j, Ano49n]. 19 [Ano38n, Ano38x, Ano39e, Ano40r, Ano41r, Ano41x, Ano41-29, Ano42l, Ano43-34, Ano44e, Ano44l, Ano44n, Ano45j, Ano45z, Ano46n, Ano47d, Ano47j, Ano47p, Ano47t, Ano48r, Ano49h].

2 [Ano38m, Ano38s, Ano38w, Ano40z, Ano41n, Ano42n, Ano44d, Ano44k, Ano45u, Ano46l, Ano46v, Ano47f, Ano49g, Ano49m]. 20 [Ano38g, Ano38p, Ano39h, Ano39x, Ano40s, Ano40w, Ano41o, Ano42w, Ano43-28, Ano43-39, Ano44h, Ano44x, Ano45q, Ano46i, Ano46o, Ano46s, Ano47g, Ano48v, Ano49c, Ano49j]. 21 [Ano38j, Ano38z, Ano39o, Ano40j, Ano40m, Ano41z, Ano42q, Ano42-27, Ano43v, Ano43-30, Ano44o, Ano45l, Ano45r, Ano45w, Ano46f, Ano47r, Ano48e, Ano48l, Ano48n, Ano49p]. 22 [Ano42f]. 25 [Ano41g]. 26th [Ano43n, Ano43q]. 27 [Ano40g]. 27th [Ano43o, Ano43t]. 28 [Ano41l]. 29 [Ano42h, Ano46c]. 29th [Ano43m, Ano43s].

3 [Ano39s, Ano40f, Ano40i, Ano40p, Ano41p, Ano43-27, Ano43-33, Ano43-37, Ano44m, Ano44s, Ano45o, Ano45y, Ano46e, Ano47i, Ano48q, Ano48u]. 30th [Ano43k, Ano43r]. 31 [Ano41k].

4 [Ano38u, Ano39m, Ano39w, Ano40l, Ano42p, Ano42v, Ano44w, Ano45h, Ano46h, Ano48k, Ano49o].

5 [Ano38o, Ano38y, Ano39f, Ano41s, Ano41y, Ano41-28, Ano42g, Ano42y, Ano43-35, Ano44f, Ano45k, Ano45f, Ano47k, Ano47q, Ano47u, Ano48s, Ano49i].

6 [Ano38h, Ano39i, Ano40n, Ano40t, Ano40x, Ano42x, Ano43-29, Ano43-38, Ano44i, Ano46j, Ano46p, Ano46t, Ano48w, Ano49d].

7 [Ano38k, Ano41-27, Ano42r, Ano42-28, Ano43u, Ano45m, Ano45s, Ano45v, Ano47s, Ano48f, Ano48m].

9 [Ano41j].

A.M. [Ano43m]. Accessibility [Ano38a, Ano38b, Ano38c, Ano38d, Ano38e, Ano39a, Ano39b, Ano39c, Ano39d, Ano40a, Ano40b, Ano40c, Ano40d, Ano40e, Ano41a, Ano41b, Ano41c, Ano41d, Ano41e, Ano41f, Ano42a, Ano42b, Ano42c, Ano42d, Ano43a, Ano43b, Ano43c, Ano43d, Ano44a, Ano44b, Ano44c, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano47a, Ano47b, Ano47c, Ano48a, Ano48b, Ano49a, Ano49b]. Additions [Hal47]. Adjourned [Ano40g, Ano40f, Ano41g, Ano42e, Ano42g, Ano42f, Ano43e]. Adopted [Ano43f]. Anniversary [Ano43g]. Appendix [Ano43f]. April [Ano38h, Ano38g, Ano39f, Ano39e, Ano40i, Ano40h, Ano41n, Ano41m, Ano42j, Ano42k, Ano42h, Ano43u, Ano43v, Ano44d, Ano44f, Ano44e, Ano45h, Ano45g, Ano46e, Ano46d, Ano47f, Ano47e, Ano48f, Ano48e, Ano49d, Ano49c]. Arrangements [Ano43f]. August [Ano38i, Ano39g, Ano40j, Ano41j, Ano41o, Ano42l, Ano43w, Ano44g, Ano45i, Ano46f, Ano47g, Ano48g, Ano49e].

Celebration [Ano43g]. Committee [Ano43f]. Corrections [Hal47].

Dec. [Ano41p, Ano41k, Ano43x, Ano44i, Ano44h]. December [Ano38k, Ano38j, Ano39i, Ano39h, Ano40l, Ano40k, Ano41q, Ano42n, Ano42m, Ano43y, Ano45k, Ano45d, Ano45j, Ano46h, Ano46g, Ano47i, Ano47h, Ano48h, Ano48i].

Eighth [Ano43k]. Errata [Ano43h]. Erratum [Ano47d].

February [Ano38m, Ano38l, Ano39j, Ano39k, Ano40n, Ano40m, Ano41s, Ano41r, Ano42p, Ano42o, Ano43-27, Ano43l, Ano43z, Ano44k, Ano44j, Ano45m, Ano45e, Ano45l, Ano46j, Ano46i, Ano47k, Ano47j, Ano48k, Ano48j, Ano49g, Ano49f]. Fifth [Ano43m]. First [Ano43n]. Fourth [Ano43o]. Front [Ano38f, Ano41h, Ano43i, Ano43j, Ano48c].

Half [Ano43k, Ano43o, Ano43q, Ano43s]. Hundredth [Ano43g].

Information [Ano40-28, Ano43-41, Ano43-42, Ano47y].

Jan. [Ano44l]. January [Ano38o, Ano38n, Ano39m, Ano39l, Ano40p, Ano40o, Ano41t, Ano41u, Ano42r, Ano42i, Ano42q, Ano43-29, Ano43-28, Ano44m, Ano45o, Ano45n, Ano46l, Ano46k, Ano47l, Ano47m, Ano48m, Ano48l, Ano49i, Ano49h]. July [Ano38p, Ano39n, Ano40f, Ano40q, Ano41v, Ano41l, Ano42e, Ano42g, Ano42f, Ano42s, Ano43-30, Ano44n, Ano45p, Ano46m, Ano47n, Ano48n, Ano49j]. June [Ano43f, Ano47d, Ano38q, Ano39o, Ano40g, Ano40r, Ano41g, Ano41w, Ano42t, Ano43-31, Ano44o, Ano45q, Ano46n, Ano47o, Ano48o, Ano49k].

Longicornia [Hal47].

M. [Ano43k, Ano43n, Ano43o, Ano43q, Ano43s, Ano43r, Ano43t]. March [Ano38s, Ano38r, Ano39p, Ano39q, Ano40t, Ano40s, Ano41y, Ano41x, Ano42v, Ano42u, Ano43-33, Ano43-32, Ano44p, Ano44q, Ano45s, Ano45r, Ano46p, Ano46o, Ano47q, Ano47p, Ano48q, Ano48p, Ano49m, Ano49l]. Matter [Ano38f, Ano41h, Ano43i, Ano43j, Ano48c]. May [Ano41i, Ano43g, Ano43k, Ano43n, Ano43o, Ano43m, Ano43q, Ano43s, Ano43r, Ano43t, Ano38u, Ano38t, Ano39s, Ano39r, Ano40u, Ano40v, Ano41-27, Ano41z, Ano42x, Ano42w, Ano43-35, Ano43p, Ano43-34, Ano44s, Ano44r, Ano45u, Ano45t, Ano46q, Ano46r, Ano46c, Ano47s, Ano47r, Ano48s, Ano48r, Ano49o, Ano49n]. Meeting [Ano38h, Ano38g, Ano38i, Ano38k, Ano38j, Ano38m, Ano38l, Ano38o, Ano38n, Ano38q, Ano38p, Ano38s, Ano38r, Ano38u, Ano38t, Ano38w, Ano38v, Ano38y, Ano38x, Ano38z, Ano39f, Ano39e, Ano39g, Ano39i, Ano39h, Ano39j, Ano39k, Ano39m, Ano39l, Ano39o, Ano39n, Ano39p, Ano39q, Ano39s, Ano39r, Ano39t, Ano39u, Ano39w, Ano39v, Ano39x, Ano40g, Ano40f, Ano40i, Ano40h, Ano40j, Ano40l, Ano40k, Ano40n, Ano40m, Ano40p, Ano40o, Ano40r, Ano40q, Ano40t, Ano40s, Ano40u, Ano40v, Ano40x, Ano40w, Ano40z, Ano40y, Ano40-27, Ano41g, Ano41i, Ano41n, Ano41m, Ano41j, Ano41o, Ano41p, Ano41q, Ano41k, Ano41s, Ano41r, Ano41t, Ano41u, Ano41w, Ano41v, Ano41l, Ano41y, Ano41x, Ano41-27, Ano41z, Ano41-28, Ano41-29]. Meeting [Ano41-30, Ano41-31, Ano41-32, Ano42e, Ano42g, Ano42f, Ano42j, Ano42k, Ano42h, Ano42l, Ano42n, Ano42m, Ano42p, Ano42o, Ano42r, Ano42i, Ano42q, Ano42t, Ano42s, Ano42v, Ano42u, Ano42x, Ano42w, Ano42y, Ano42z, Ano42-28, Ano42-27, Ano42-29, Ano43e, Ano43u, Ano43v, Ano43w, Ano43y, Ano43x, Ano43k, Ano43-27, Ano43l, Ano43z, Ano43n, Ano43o, Ano43m, Ano43-29, Ano43-28, Ano43-31, Ano43-30, Ano43-33, Ano43-32, Ano43-35, Ano43p, Ano43-34, Ano43-37, Ano43-36, Ano43-38, Ano43-39, Ano43q, Ano43s, Ano43r, Ano43-40, Ano43t, Ano44d, Ano44f, Ano44e, Ano44g, Ano44i, Ano44h, Ano44k, Ano44j, Ano44m, Ano44l, Ano44o, Ano44n, Ano44p, Ano44q, Ano44s, Ano44r, Ano44t, Ano44u, Ano44w, Ano44v, Ano44x, Ano45h, Ano45g, Ano45i, Ano45k]. Meeting [Ano45d, Ano45j, Ano45m, Ano45e, Ano45l, Ano45o, Ano45n, Ano45q, Ano45p, Ano45s, Ano45r, Ano45u, Ano45t, Ano45v, Ano45w, Ano45y, Ano45x, Ano45f, Ano45z, Ano46e, Ano46d, Ano46f, Ano46h, Ano46g, Ano46j, Ano46i, Ano46l, Ano46k, Ano46n, Ano46m, Ano46p, Ano46o, Ano46q, Ano46r, Ano46c, Ano46t, Ano46s, Ano46v, Ano46u, Ano46w, Ano47d, Ano47f, Ano47e, Ano47g, Ano47i, Ano47h, Ano47k, Ano47j, Ano47l, Ano47m, Ano47o, Ano47n, Ano47q, Ano47p, Ano47s, Ano47r, Ano47u, Ano47t, Ano47v, Ano47w, Ano47x, Ano48f, Ano48e, Ano48g, Ano48h, Ano48i, Ano48k, Ano48j, Ano48m, Ano48l, Ano48o, Ano48n, Ano48q, Ano48p, Ano48s, Ano48r, Ano48u, Ano48t, Ano48w, Ano48d, Ano48v, Ano48x, Ano49d, Ano49c]. Meeting [Ano49e, Ano49g, Ano49f, Ano49i, Ano49h, Ano49k, Ano49j, Ano49m, Ano49l, Ano49o, Ano49n, Ano49p].

Note [Ano41i]. Nov. [Ano41-28, Ano42y, Ano44t, Ano45v]. November [Ano38w, Ano38v, Ano39t, Ano39u, Ano40x, Ano40w, Ano41-29, Ano42z, Ano43-37, Ano43-36, Ano44u, Ano45w, Ano46t, Ano46s, Ano47u, Ano47t, Ano48u, Ano48t].

O’Clock [Ano43k, Ano43n, Ano43o, Ano43m, Ano43q, Ano43s, Ano43r, Ano43t]. Oct.

[Ano42-27, Ano43e, Ano43-38]. October [Ano38y, Ano38x, Ano39w, Ano39v, Ano40z, Ano40y, Ano41-30, Ano41-31, Ano42-28, Ano43-39, Ano44w, Ano44v, Ano45y, Ano45x, Ano46v, Ano46u, Ano47v, Ano47w, Ano48w, Ano48d, Ano48v].

P [Ano43k, Ano43o, Ano43q, Ano43s]. Paper [Hal47]. Past [Ano43k, Ano43o, Ano43q, Ano43s]. Presented [Ano43f].

Report [Ano43f]. Resolutions [Ano43f].

Second [Ano43q]. September [Ano38z, Ano39x, Ano40-27, Ano41-32, Ano42-29, Ano43-40, Ano44x, Ano45f, Ano45z, Ano46w, Ano47x, Ano48x, Ano49p]. Session [Ano43k, Ano43n, Ano43o, Ano43m, Ano43q, Ano43s, Ano43r, Ano43t]. Seven [Ano43q]. Seventh [Ano43r]. Sixth [Ano43s]. Society [Ano43f]. Special [Ano41j, Ano41k, Ano41l, Ano42h, Ano42i, Ano43k, Ano43l, Ano43n, Ano43o, Ano43m, Ano43p, Ano43q, Ano43s, Ano43r, Ano43t, Ano44d, Ano45d, Ano45e, Ano45f, Ano46c, Ano48d]. State [Ano43u]. Stated [Ano38h, Ano38g, Ano38i, Ano38k, Ano38j, Ano38m, Ano38l, Ano38o, Ano38n, Ano38q, Ano38p, Ano38s, Ano38r, Ano38u, Ano38t, Ano38w, Ano38v, Ano38y, Ano38x, Ano38z, Ano39f, Ano39e, Ano39g, Ano39i, Ano39h, Ano39j, Ano39k, Ano39m, Ano39l, Ano39o, Ano39n, Ano39p, Ano39q, Ano39s, Ano39r, Ano39t, Ano39u, Ano39w, Ano39v, Ano39x, Ano40i, Ano40h, Ano40j, Ano40l, Ano40k, Ano40n, Ano40m, Ano40p, Ano40o, Ano40r, Ano40q, Ano40t, Ano40s, Ano40u, Ano40v, Ano40x, Ano40w, Ano40z, Ano40y, Ano40-27, Ano41i, Ano41n, Ano41m, Ano41o, Ano41p, Ano41q, Ano41s, Ano41r, Ano41t, Ano41u, Ano41w, Ano41v, Ano41y, Ano41x, Ano41-27, Ano41z, Ano41-28, Ano41-29, Ano41-30, Ano41-31, Ano41-32, Ano42j, Ano42k, Ano42l]. Stated [Ano42n, Ano42m, Ano42p, Ano42o, Ano42r, Ano42q, Ano42t, Ano42s, Ano42v, Ano42u, Ano42x, Ano42w, Ano42y, Ano42z, Ano42-28, Ano42-27, Ano42-29, Ano43v, Ano43w, Ano43y, Ano43x, Ano43-27, Ano43z, Ano43-29, Ano43-28, Ano43-31, Ano43-30, Ano43-33, Ano43-32, Ano43-35, Ano43-34, Ano43-37, Ano43-36, Ano43-38, Ano43-39, Ano43-40, Ano44f, Ano44e, Ano44g, Ano44i, Ano44h, Ano44k, Ano44j, Ano44m, Ano44l, Ano44o, Ano44n, Ano44p, Ano44q, Ano44s, Ano44r, Ano44t, Ano44u, Ano44w, Ano44v, Ano44x, Ano45h, Ano45g, Ano45i, Ano45k, Ano45j, Ano45m, Ano45l, Ano45o, Ano45n, Ano45q, Ano45p, Ano45s, Ano45r, Ano45u, Ano45t, Ano45v, Ano45w, Ano45y, Ano45x, Ano45z, Ano46e, Ano46d, Ano46f, Ano46h, Ano46g, Ano46j, Ano46i, Ano46l]. Stated [Ano46k, Ano46n, Ano46m, Ano46p, Ano46o, Ano46q, Ano46r, Ano46t, Ano46s, Ano46v, Ano46u, Ano46w, Ano47f, Ano47e, Ano47g, Ano47i, Ano47h, Ano47k, Ano47j, Ano47l, Ano47m, Ano47o, Ano47n, Ano47q, Ano47p, Ano47s, Ano47r, Ano47u, Ano47t, Ano47v, Ano47w, Ano47x, Ano48f, Ano48e, Ano48g, Ano48h, Ano48i, Ano48k, Ano48j, Ano48m, Ano48l, Ano48o, Ano48n, Ano48q, Ano48p, Ano48s, Ano48r, Ano48u, Ano48t, Ano48w, Ano48v, Ano48x, Ano49d, Ano49c, Ano49e, Ano49g, Ano49f, Ano49i, Ano49h, Ano49k, Ano49j, Ano49m, Ano49l, Ano49o, Ano49n, Ano49p, Ano47d]. States [Hal47].

Thereupon [Ano43f]. Third [Ano43t].

United [Hal47].

Volume [Ano40-28, Ano43-41, Ano43-42, Ano47y].

References

Anonymous:1838:Aa

[Ano38a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(1):??, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1838:Ab

[Ano38b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(2):??, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1838:Ac

[Ano38c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(3):??, July/August 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1838:Ad

[Ano38d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(4):??, September/October 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1838:Ae

[Ano38e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(5):??, November/December 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1838:FM

[Ano38f] Anonymous. Front matter. Proceedings of the American Philosophical Society, 1(1):??, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

          www.jstor.org/stable/3143185.

Anonymous:1838:SMAb

[Ano38g] Anonymous. Stated meeting, April 20. Proceedings of the American Philosophical Society, 1(2):11, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143202.

Anonymous:1838:SMAa

[Ano38h] Anonymous. Stated meeting, April 6. Proceedings of the American Philosophical Society, 1(2):9–10, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143201.

Anonymous:1838:SMAc

[Ano38i] Anonymous. Stated meeting, August 17. Proceedings of the American Philosophical Society, 1(3):29–31, July/August 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143503.

Anonymous:1838:SMDb

[Ano38j] Anonymous. Stated meeting, December 21. Proceedings of the American Philosophical Society, 1(5):62–66, November/December 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143099.

Anonymous:1838:SMDa

[Ano38k] Anonymous. Stated meeting, December 7. Proceedings of the American Philosophical Society, 1(5):60–62, November/December 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143098.

Anonymous:1838:SMFb

[Ano38l] Anonymous. Stated meeting, February 16. Proceedings of the American Philosophical Society, 1(1):5–6, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143189.

Anonymous:1838:SMFa

[Ano38m] Anonymous. Stated meeting, February 2. Proceedings of the American Philosophical Society, 1(1):4–5, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143188.

Anonymous:1838:SMJb

[Ano38n] Anonymous. Stated meeting, January 19. Proceedings of the American Philosophical Society, 1(1):2–3, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143187.

Anonymous:1838:SMJa

[Ano38o] Anonymous. Stated meeting, January 5. Proceedings of the American Philosophical Society, 1(1):1–2, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143186.

Anonymous:1838:SMJd

[Ano38p] Anonymous. Stated meeting, July 20. Proceedings of the American Philosophical Society, 1(3):21–28, July/August 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143502.

Anonymous:1838:SMJc

[Ano38q] Anonymous. Stated meeting, June 15. Proceedings of the American Philosophical Society, 1(2):16–19, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143205.

Anonymous:1838:SMMb

[Ano38r] Anonymous. Stated meeting, March 16. Proceedings of the American Philosophical Society, 1(1):8, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143191.

Anonymous:1838:SMMa

[Ano38s] Anonymous. Stated meeting, March 2. Proceedings of the American Philosophical Society, 1(1):6–7, January/March 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143190.

Anonymous:1838:SMMd

[Ano38t] Anonymous. Stated meeting, May 18. Proceedings of the American Philosophical Society, 1(2):15–16, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143204.

Anonymous:1838:SMMc

[Ano38u] Anonymous. Stated meeting, May 4. Proceedings of the American Philosophical Society, 1(2):12–15, April/June 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143203.

Anonymous:1838:SMNb

[Ano38v] Anonymous. Stated meeting, November 16. Proceedings of the American Philosophical Society, 1(5):56–60, November/December 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143097.

Anonymous:1838:SMNa

[Ano38w] Anonymous. Stated meeting, November 2. Proceedings of the American Philosophical Society, 1(5):49–56, November/December 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143096.

Anonymous:1838:SMOb

[Ano38x] Anonymous. Stated meeting, October 19. Proceedings of the American Philosophical Society, 1(4):47–48, September/October 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143453.

Anonymous:1838:SMOa

[Ano38y] Anonymous. Stated meeting, October 5. Proceedings of the American Philosophical Society, 1(4):42–47, September/October 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143452.

Anonymous:1838:SMS

[Ano38z] Anonymous. Stated meeting, September 21. Proceedings of the American Philosophical Society, 1(4):33–42, September/October 1838. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143451.

Anonymous:1839:Aa

[Ano39a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(6):??, January/March 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1839:Ab

[Ano39b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(7):??, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1839:Ac

[Ano39c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(8):??, September/October 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1839:Ad

[Ano39d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(9):??, November/December 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1839:SMAb

[Ano39e] Anonymous. Stated meeting, April 19. Proceedings of the American Philosophical Society, 1(7):90–95, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143296.

Anonymous:1839:SMAa

[Ano39f] Anonymous. Stated meeting, April 5. Proceedings of the American Philosophical Society, 1(7):87–90, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143295.

Anonymous:1839:SMAc

[Ano39g] Anonymous. Stated meeting, August 16. Proceedings of the American Philosophical Society, 1(7):109–111, April/August 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143301.

Anonymous:1839:SMDb

[Ano39h] Anonymous. Stated meeting, December 20. Proceedings of the American Philosophical Society, 1(9):157–163, November/December 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143278.

Anonymous:1839:SMDa

[Ano39i] Anonymous. Stated meeting, December 6. Proceedings of the American Philosophical Society, 1(9):149–157, November/December 1839.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143277.

Anonymous:1839:SMFa

[Ano39j] Anonymous. Stated meeting, February 1. Proceedings of the American Philosophical Society, 1(6):73–74, January/March 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143113.

Anonymous:1839:SMFb

[Ano39k] Anonymous. Stated meeting, February 15. Proceedings of the American Philosophical Society, 1(6):75–78, January/March 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143114.

Anonymous:1839:SMJb

[Ano39l] Anonymous. Stated meeting, January 18. Proceedings of the American Philosophical Society, 1(6):70–73, January/March 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143112.

Anonymous:1839:SMJa

[Ano39m] Anonymous. Stated meeting, January 4. Proceedings of the American Philosophical Society, 1(6):67–69, January/March 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143111.

Anonymous:1839:SMJd

[Ano39n] Anonymous. Stated meeting, July 17. Proceedings of the American Philosophical Society, 1(7):105–108, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143300.

Anonymous:1839:SMJc

[Ano39o] Anonymous. Stated meeting, June 21. Proceedings of the American Philosophical Society, 1(7):102–105, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143299.

Anonymous:1839:SMMa

[Ano39p] Anonymous. Stated meeting, March 1. Proceedings of the American Philosophical Society, 1(6):78–80, January/March 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143115.

Anonymous:1839:SMMb

[Ano39q] Anonymous. Stated meeting, March 15. Proceedings of the American Philosophical Society, 1(6):81–85, January/March 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143116.

Anonymous:1839:SMMd

[Ano39r] Anonymous. Stated meeting, May 17. Proceedings of the American Philosophical Society, 1(7):98–102, April/August 1839. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143298.

Anonymous:1839:SMMc

[Ano39s] Anonymous. Stated meeting, May 3. Proceedings of the American Philosophical Society, 1(7):95–98, April/August 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143297.

Anonymous:1839:SMNa

[Ano39t] Anonymous. Stated meeting, November 1. Proceedings of the American Philosophical Society, 1(9):137–139, November/December 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143275.

Anonymous:1839:SMNb

[Ano39u] Anonymous. Stated meeting, November 15. Proceedings of the American Philosophical Society, 1(9):140–149, November/December 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143276.

Anonymous:1839:SMOb

[Ano39v] Anonymous. Stated meeting, October 18. Proceedings of the American Philosophical Society, 1(8):126–136, September/October 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143290.

Anonymous:1839:SMOa

[Ano39w] Anonymous. Stated meeting, October 4. Proceedings of the American Philosophical Society, 1(8):118–126, September/October 1839.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143289.

Anonymous:1839:SMS

[Ano39x] Anonymous. Stated meeting, September 20. Proceedings of the American Philosophical Society, 1(8):113–118, September/October 1839. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143288.

Anonymous:1840:Aa

[Ano40a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(10):??, January/February 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1840:Ab

[Ano40b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(11):??, March/April 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1840:Ac

[Ano40c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(12):??, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1840:Ad

[Ano40d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(13):??, August/October 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1840:Ae

[Ano40e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 1(14):??, November/December 1840. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1840:AMJb

[Ano40f] Anonymous. Adjourned meeting, July 3. Proceedings of the American Philosophical Society, 1(12):234–243, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143109.

Anonymous:1840:AMJa

[Ano40g] Anonymous. Adjourned meeting, June 27. Proceedings of the American Philosophical Society, 1(12):234, May/July 1840. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143108.

Anonymous:1840:SMAb

[Ano40h] Anonymous. Stated meeting, April 17. Proceedings of the American Philosophical Society, 1(11):194–201, March/April 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143257.

Anonymous:1840:SMAa

[Ano40i] Anonymous. Stated meeting, April 3. Proceedings of the American Philosophical Society, 1(11):189–193, March/April 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143256.

Anonymous:1840:SMAc

[Ano40j] Anonymous. Stated meeting, August 21. Proceedings of the American Philosophical Society, 1(13):257–267, August/October 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143528.

Anonymous:1840:SMDb

[Ano40k] Anonymous. Stated meeting, December 18. Proceedings of the American Philosophical Society, 1(14):313–325, November/ December 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143483.

Anonymous:1840:SMDa

[Ano40l] Anonymous. Stated meeting, December 4. Proceedings of the American Philosophical Society, 1(14):304–313, November/December 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143482.

Anonymous:1840:SMFb

[Ano40m] Anonymous. Stated meeting, February 21. Proceedings of the American Philosophical Society, 1(10):172–178, January/February 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143294.

Anonymous:1840:SMFa

[Ano40n] Anonymous. Stated meeting, February 6. Proceedings of the American Philosophical Society, 1(10):170–171, January/February 1840.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143293.

Anonymous:1840:SMJb

[Ano40o] Anonymous. Stated meeting, January 17. Proceedings of the American Philosophical Society, 1(10):167–169, January/February 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143292.

Anonymous:1840:SMJa

[Ano40p] Anonymous. Stated meeting, January 3. Proceedings of the American Philosophical Society, 1(10):165–167, January/February 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143291.

Anonymous:1840:SMJd

[Ano40q] Anonymous. Stated meeting, July 17. Proceedings of the American Philosophical Society, 1(12):243–256, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143110.

Anonymous:1840:SMJc

[Ano40r] Anonymous. Stated meeting, June 19. Proceedings of the American Philosophical Society, 1(12):219–233, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143107.

Anonymous:1840:SMMb

[Ano40s] Anonymous. Stated meeting, March 20. Proceedings of the American Philosophical Society, 1(11):182–188, March/April 1840. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143255.

Anonymous:1840:SMMa

[Ano40t] Anonymous. Stated meeting, March 6. Proceedings of the American Philosophical Society, 1(11):179–181, March/April 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143254.

Anonymous:1840:SMMc

[Ano40u] Anonymous. Stated meeting, May 1. Proceedings of the American Philosophical Society, 1(12):203–209, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143105.

Anonymous:1840:SMMd

[Ano40v] Anonymous. Stated meeting, May 15. Proceedings of the American Philosophical Society, 1(12):209–218, May/July 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143106.

Anonymous:1840:SMNb

[Ano40w] Anonymous. Stated meeting, November 20. Proceedings of the American Philosophical Society, 1(14):296–304, November/ December 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143481.

Anonymous:1840:SMNa

[Ano40x] Anonymous. Stated meeting, November 6. Proceedings of the American Philosophical Society, 1(14):291–295, November/December 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143480.

Anonymous:1840:SMOb

[Ano40y] Anonymous. Stated meeting, October 16. Proceedings of the American Philosophical Society, 1(13):283–290, August/October 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143531.

Anonymous:1840:SMOa

[Ano40z] Anonymous. Stated meeting, October 2. Proceedings of the American Philosophical Society, 1(13):276–283, August/October 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143530.

Anonymous:1840:SMS

[Ano40-27] Anonymous. Stated meeting, September 18. Proceedings of the American Philosophical Society, 1(13):267–276, August/October 1840. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143529.

Anonymous:1840:VI

[Ano40-28] Anonymous. Volume information. Proceedings of the American Philosophical Society, 1(14):327–336, November/December 1840.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143479.

Anonymous:1841:Aa

[Ano41a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(15):??, January 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:Ab

[Ano41b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(16):??, February 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:Ac

[Ano41c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(17):??, March/April 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:Ad

[Ano41d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(18):??, May/June 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:Ae

[Ano41e] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(19):??, July/October 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:Af

[Ano41f] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(20):??, November/December 1841. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1841:AMJ

[Ano41g] Anonymous. Adjourned meeting, June 25. Proceedings of the American Philosophical Society, 2(18):77–78, May/June 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143095.

Anonymous:1841:FM

[Ano41h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 2(15):??, January 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/3143448.

Anonymous:1841:NSM

[Ano41i] Anonymous. Note: Stated meeting, may 21. Proceedings of the American Philosophical Society, 2(19):106, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143265.

Anonymous:1841:SMAc

[Ano41j] Anonymous. Special meeting, August 9. Proceedings of the American Philosophical Society, 2(19):86–87, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143260.

Anonymous:1841:SMDc

[Ano41k] Anonymous. Special meeting, Dec. 31. Proceedings of the American Philosophical Society, 2(20):131–133, November/December 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143104.

Anonymous:1841:SMJe

[Ano41l] Anonymous. Special meeting, July 28. Proceedings of the American Philosophical Society, 2(19):86, July/October 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143259.

Anonymous:1841:SMAb

[Ano41m] Anonymous. Stated meeting, April 16. Proceedings of the American Philosophical Society, 2(17):43–46, March/April 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143081.

Anonymous:1841:SMAa

[Ano41n] Anonymous. Stated meeting, April 2. Proceedings of the American Philosophical Society, 2(17):39–43, March/April 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143080.

Anonymous:1841:SMAd

[Ano41o] Anonymous. Stated meeting, August 20. Proceedings of the American Philosophical Society, 2(19):87–93, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143261.

Anonymous:1841:SMDa

[Ano41p] Anonymous. Stated meeting, Dec. 3. Proceedings of the American Philosophical Society, 2(20):119–125, November/December 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143102.

Anonymous:1841:SMDb

[Ano41q] Anonymous. Stated meeting, December 17. Proceedings of the American Philosophical Society, 2(20):125–130, November/ December 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143103.

Anonymous:1841:SMFb

[Ano41r] Anonymous. Stated meeting, February 19. Proceedings of the American Philosophical Society, 2(16):17–28, February 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143193.

Anonymous:1841:SMFa

[Ano41s] Anonymous. Stated meeting, February 5. Proceedings of the American Philosophical Society, 2(16):9–16, February 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143192.

Anonymous:1841:SMJa

[Ano41t] Anonymous. Stated meeting, January 1. Proceedings of the American Philosophical Society, 2(15):1–4, January 1841. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143449.

Anonymous:1841:SMJb

[Ano41u] Anonymous. Stated meeting, January 15. Proceedings of the American Philosophical Society, 2(15):4–8, January 1841. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143450.

Anonymous:1841:SMJd

[Ano41v] Anonymous. Stated meeting, July 16. Proceedings of the American Philosophical Society, 2(19):79–86, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143258.

Anonymous:1841:SMJc

[Ano41w] Anonymous. Stated meeting, June 18. Proceedings of the American Philosophical Society, 2(18):70–77, May/June 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143094.

Anonymous:1841:SMMb

[Ano41x] Anonymous. Stated meeting, March 19. Proceedings of the American Philosophical Society, 2(17):36–38, March/April 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143079.

Anonymous:1841:SMMa

[Ano41y] Anonymous. Stated meeting, March 5. Proceedings of the American Philosophical Society, 2(17):29–36, March/April 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143078.

Anonymous:1841:SMMd

[Ano41z] Anonymous. Stated meeting, May 21. Proceedings of the American Philosophical Society, 2(18):58–70, May/June 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143093.

Anonymous:1841:SMMc

[Ano41-27] Anonymous. Stated meeting, May 7. Proceedings of the American Philosophical Society, 2(18):47–58, May/June 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143092.

Anonymous:1841:SMNa

[Ano41-28] Anonymous. Stated meeting, Nov. 5. Proceedings of the American Philosophical Society, 2(20):107–117, November/December 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143100.

Anonymous:1841:SMNb

[Ano41-29] Anonymous. Stated meeting, November 19. Proceedings of the American Philosophical Society, 2(20):117–119, November/ December 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143101.

Anonymous:1841:SMOa

[Ano41-30] Anonymous. Stated meeting, October 1. Proceedings of the American Philosophical Society, 2(19):99–103, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143263.

Anonymous:1841:SMOb

[Ano41-31] Anonymous. Stated meeting, October 15. Proceedings of the American Philosophical Society, 2(19):103–106, July/October 1841. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143264.

Anonymous:1841:SMS

[Ano41-32] Anonymous. Stated meeting, September 17. Proceedings of the American Philosophical Society, 2(19):93–99, July/October 1841. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143262.

Anonymous:1842:Aa

[Ano42a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(21):??, January/April 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1842:Ab

[Ano42b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(22):??, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1842:Ac

[Ano42c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(23):??, August/October 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1842:Ad

[Ano42d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(24):??, November/December 1842. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1842:AMJa

[Ano42e] Anonymous. Adjourned meeting, July 1. Proceedings of the American Philosophical Society, 2(22):197, May/July 1842. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143548.

Anonymous:1842:AMJc

[Ano42f] Anonymous. Adjourned meeting, July 22. Proceedings of the American Philosophical Society, 2(22):205–206, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143551.

Anonymous:1842:AMJb

[Ano42g] Anonymous. Adjourned meeting, July 5. Proceedings of the American Philosophical Society, 2(22):197–198, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143549.

Anonymous:1842:SMAc

[Ano42h] Anonymous. Special meeting, April 29. Proceedings of the American Philosophical Society, 2(21):171, January/April 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143184.

Anonymous:1842:SMJb

[Ano42i] Anonymous. Special meeting, January 15. Proceedings of the American Philosophical Society, 2(21):138–139, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143176.

Anonymous:1842:SMAa

[Ano42j] Anonymous. Stated meeting, April 1. Proceedings of the American Philosophical Society, 2(21):158–162, January/April 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143182.

Anonymous:1842:SMAb

[Ano42k] Anonymous. Stated meeting, April 15. Proceedings of the American Philosophical Society, 2(21):162–171, January/April 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143183.

Anonymous:1842:SMAd

[Ano42l] Anonymous. Stated meeting, August 19. Proceedings of the American Philosophical Society, 2(23):207–210, August/October 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143415.

Anonymous:1842:SMDb

[Ano42m] Anonymous. Stated meeting, December 16. Proceedings of the American Philosophical Society, 2(24):238–243, November/ December 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143305.

Anonymous:1842:SMDa

[Ano42n] Anonymous. Stated meeting, December 2. Proceedings of the American Philosophical Society, 2(24):236–238, November/December 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143304.

Anonymous:1842:SMFb

[Ano42o] Anonymous. Stated meeting, February 18. Proceedings of the American Philosophical Society, 2(21):148–149, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143179.

Anonymous:1842:SMFa

[Ano42p] Anonymous. Stated meeting, February 4. Proceedings of the American Philosophical Society, 2(21):145–148, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143178.

Anonymous:1842:SMJc

[Ano42q] Anonymous. Stated meeting, January 21. Proceedings of the American Philosophical Society, 2(21):140–145, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143177.

Anonymous:1842:SMJa

[Ano42r] Anonymous. Stated meeting, January 7. Proceedings of the American Philosophical Society, 2(21):135–138, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143175.

Anonymous:1842:SMJe

[Ano42s] Anonymous. Stated meeting, July 15. Proceedings of the American Philosophical Society, 2(22):198–205, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143550.

Anonymous:1842:SMJd

[Ano42t] Anonymous. Stated meeting, June 17. Proceedings of the American Philosophical Society, 2(22):191–197, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143547.

Anonymous:1842:SMMb

[Ano42u] Anonymous. Stated meeting, March 18. Proceedings of the American Philosophical Society, 2(21):151–157, January/April 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143181.

Anonymous:1842:SMMa

[Ano42v] Anonymous. Stated meeting, March 4. Proceedings of the American Philosophical Society, 2(21):149–151, January/April 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143180.

Anonymous:1842:SMMd

[Ano42w] Anonymous. Stated meeting, May 20. Proceedings of the American Philosophical Society, 2(22):184–191, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143546.

Anonymous:1842:SMMc

[Ano42x] Anonymous. Stated meeting, May 6. Proceedings of the American Philosophical Society, 2(22):173–184, May/July 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143545.

Anonymous:1842:SMNa

[Ano42y] Anonymous. Stated meeting, Nov. 5. Proceedings of the American Philosophical Society, 2(24):231–233, November/December 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143302.

Anonymous:1842:SMNb

[Ano42z] Anonymous. Stated meeting, November 18. Proceedings of the American Philosophical Society, 2(24):234–236, November/ December 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/3143303.

Anonymous:1842:SMOb

[Ano42-27] Anonymous. Stated meeting, Oct. 21. Proceedings of the American Philosophical Society, 2(23):222–230, August/October 1842. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143418.

Anonymous:1842:SMOa

[Ano42-28] Anonymous. Stated meeting, October 7. Proceedings of the American Philosophical Society, 2(23):214–222, August/October 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143417.

Anonymous:1842:SMS

[Ano42-29] Anonymous. Stated meeting, September 16. Proceedings of the American Philosophical Society, 2(23):211–214, August/October 1842. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143416.

Anonymous:1843:Aa

[Ano43a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(25):??, January/March 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1843:Ab

[Ano43b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 2(26):??, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1843:Ac

[Ano43c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 3(27):??, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1843:Ad

[Ano43d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(28):??, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1843:AMO

[Ano43e] Anonymous. Adjourned meeting, Oct. 13. Proceedings of the American Philosophical Society, 4(28):17–18, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981577.

Anonymous:1843:ARC

[Ano43f] Anonymous. Appendix. Report of the Committee of Arrangements, Presented June 16, 1843, and Resolutions Adopted Thereupon by the Society. Proceedings of the American Philosophical Society, 3 (27):213–218, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982022.

Anonymous:1843:CHA

[Ano43g] Anonymous. Celebration of the hundredth anniversary, may 25, 1843. Proceedings of the American Philosophical Society, 3(27):1–36, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982013.

Anonymous:1843:E

[Ano43h] Anonymous. Errata. Proceedings of the American Philosophical Society, 3(27):232, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982023.

Anonymous:1843:FMa

[Ano43i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 3(27):219–232, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/982012.

Anonymous:1843:FMb

[Ano43j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 4(28):??, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/981571.

Anonymous:1843:SME

[Ano43k] Anonymous. Special meeting. Eighth Session, 30th May, Half Past 5 O’Clock, P. M. Proceedings of the American Philosophical Society, 3(27):169–211, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982021.

Anonymous:1843:SMFb

[Ano43l] Anonymous. Special meeting, February 10. Proceedings of the American Philosophical Society, 2(25):252–253, January/March 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143149.

Anonymous:1843:SMFf

[Ano43m] Anonymous. Special meeting. Fifth Session, 29th May, 10 O’Clock, A.M. Proceedings of the American Philosophical Society, 3(27):115– 138, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982018.

Anonymous:1843:SMFd

[Ano43n] Anonymous. Special meeting. First Session, 26th May, 1843, 10 O’Clock, A. M. Proceedings of the American Philosophical Society, 3(27):37–56, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982014.

Anonymous:1843:SMFe

[Ano43o] Anonymous. Special meeting. Fourth Session, 27th May, Half Past 7 O’Clock, P. M. Proceedings of the American Philosophical Society, 3(27):102–114, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982017.

Anonymous:1843:SMMd

[Ano43p] Anonymous. Special meeting, May 12. Proceedings of the American Philosophical Society, 2(26):273, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143076.

Anonymous:1843:SMSa

[Ano43q] Anonymous. Special meeting. Second Session, 26th May, Half Past Seven O’Clock, P. M. Proceedings of the American Philosophical Society, 3(27):57–67, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982015.

Anonymous:1843:SMSc

[Ano43r] Anonymous. Special meeting. Seventh Session, 30th May, 10 O’Clock, A. M. Proceedings of the American Philosophical Society, 3(27):151–169, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982020.

Anonymous:1843:SMSb

[Ano43s] Anonymous. Special meeting. Sixth Session, 29th May, Half Past 7 O’Clock, P. M. Proceedings of the American Philosophical Society, 3(27):138–150, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/982019.

Anonymous:1843:SMT

[Ano43t] Anonymous. Special meeting. Third Session, 27th May, 10 O’Clock, A. M. Proceedings of the American Philosophical Society, 3(27):67– 102, May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982016.

Anonymous:1843:SMAa

[Ano43u] Anonymous. State meeting, April 7. Proceedings of the American Philosophical Society, 2(26):263–268, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143073.

Anonymous:1843:SMAb

[Ano43v] Anonymous. Stated meeting, April 21. Proceedings of the American Philosophical Society, 2(26):268–270, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143074.

Anonymous:1843:SMAc

[Ano43w] Anonymous. Stated meeting, August 18. Proceedings of the American Philosophical Society, 4(28):6–8, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981574.

Anonymous:1843:SMDb

[Ano43x] Anonymous. Stated meeting, Dec. 15. Proceedings of the American Philosophical Society, 4(28):29–30, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981582.

Anonymous:1843:SMDa

[Ano43y] Anonymous. Stated meeting, December 1. Proceedings of the American Philosophical Society, 4(28):26–29, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981581.

Anonymous:1843:SMFc

[Ano43z] Anonymous. Stated meeting, February 17. Proceedings of the American Philosophical Society, 2(25):253–254, January/March 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143150.

Anonymous:1843:SMFa

[Ano43-27] Anonymous. Stated meeting, February 3. Proceedings of the American Philosophical Society, 2(25):251–252, January/March 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143148.

Anonymous:1843:SMJb

[Ano43-28] Anonymous. Stated meeting, January 20. Proceedings of the American Philosophical Society, 2(25):248–251, January/March 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143147.

Anonymous:1843:SMJa

[Ano43-29] Anonymous. Stated meeting, January 6. Proceedings of the American Philosophical Society, 2(25):245–247, January/March 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143146.

Anonymous:1843:SMJd

[Ano43-30] Anonymous. Stated meeting, July 21. Proceedings of the American Philosophical Society, 4(28):4–6, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981573.

Anonymous:1843:SMJc

[Ano43-31] Anonymous. Stated meeting, June 16. Proceedings of the American Philosophical Society, 4(28):1–4, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981572.

Anonymous:1843:SMMb

[Ano43-32] Anonymous. Stated meeting, March 17. Proceedings of the American Philosophical Society, 2(25):256–262, January/March 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143152.

Anonymous:1843:SMMa

[Ano43-33] Anonymous. Stated meeting, March 3. Proceedings of the American Philosophical Society, 2(25):254–256, January/March 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143151.

Anonymous:1843:SMMe

[Ano43-34] Anonymous. Stated meeting, May 19. Proceedings of the American Philosophical Society, 2(26):274–276, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143077.

Anonymous:1843:SMMc

[Ano43-35] Anonymous. Stated meeting, May 5. Proceedings of the American Philosophical Society, 2(26):270–273, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143075.

Anonymous:1843:SMNb

[Ano43-36] Anonymous. Stated meeting, November 17. Proceedings of the American Philosophical Society, 4(28):24–26, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981580.

Anonymous:1843:SMNa

[Ano43-37] Anonymous. Stated meeting, November 3. Proceedings of the American Philosophical Society, 4(28):20–23, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981579.

Anonymous:1843:SMOa

[Ano43-38] Anonymous. Stated meeting, Oct. 6. Proceedings of the American Philosophical Society, 4(28):13–16, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981576.

Anonymous:1843:SMOb

[Ano43-39] Anonymous. Stated meeting, October 20. Proceedings of the American Philosophical Society, 4(28):18–20, June/December 1843. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981578.

Anonymous:1843:SMSd

[Ano43-40] Anonymous. Stated meeting, September 15. Proceedings of the American Philosophical Society, 4(28):8–13, June/December 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981575.

Anonymous:1843:VIa

[Ano43-41] Anonymous. Volume information. Proceedings of the American Philosophical Society, 2(26):277–282, April/May 1843. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143072.

Anonymous:1843:VIb

[Ano43-42] Anonymous. Volume information. Proceedings of the American Philosophical Society, 4(28):??, June/December 1843. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981570.

Anonymous:1844:Aa

[Ano44a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(29):??, January/April 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1844:Ab

[Ano44b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(30):??, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1844:Ac

[Ano44c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(31):??, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1844:SMAa

[Ano44d] Anonymous. Special meeting, April 2. Proceedings of the American Philosophical Society, 4(29):51, January/April 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981977.

Anonymous:1844:SMAc

[Ano44e] Anonymous. Stated meeting, April 19. Proceedings of the American Philosophical Society, 4(30):59–66, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982201.

Anonymous:1844:SMAb

[Ano44f] Anonymous. Stated meeting, April, 5. Proceedings of the American Philosophical Society, 4(30):53–59, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982200.

Anonymous:1844:SMAd

[Ano44g] Anonymous. Stated meeting, August 16th. Proceedings of the American Philosophical Society, 4(31):96–99, July/December 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982082.

Anonymous:1844:SMDb

[Ano44h] Anonymous. Stated meeting, Dec. 20. Proceedings of the American Philosophical Society, 4(31):123–131, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982089.

Anonymous:1844:SMDa

[Ano44i] Anonymous. Stated meeting, Dec. 6. Proceedings of the American Philosophical Society, 4(31):115–123, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982088.

Anonymous:1844:SMFb

[Ano44j] Anonymous. Stated meeting, February 16. Proceedings of the American Philosophical Society, 4(29):43–44, January/April 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981974.

Anonymous:1844:SMFa

[Ano44k] Anonymous. Stated meeting, February 2. Proceedings of the American Philosophical Society, 4(29):39–43, January/April 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981973.

Anonymous:1844:SMJb

[Ano44l] Anonymous. Stated meeting, Jan. 19. Proceedings of the American Philosophical Society, 4(29):37–39, January/April 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981972.

Anonymous:1844:SMJa

[Ano44m] Anonymous. Stated meeting, January 3. Proceedings of the American Philosophical Society, 4(29):31–36, January/April 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981971.

Anonymous:1844:SMJd

[Ano44n] Anonymous. Stated meeting, July 19. Proceedings of the American Philosophical Society, 4(31):93–96, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982081.

Anonymous:1844:SMJc

[Ano44o] Anonymous. Stated meeting, June 21. Proceedings of the American Philosophical Society, 4(30):86–92, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982204.

Anonymous:1844:SMMa

[Ano44p] Anonymous. Stated meeting, March 1. Proceedings of the American Philosophical Society, 4(29):44–47, January/April 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981975.

Anonymous:1844:SMMb

[Ano44q] Anonymous. Stated meeting, March 15. Proceedings of the American Philosophical Society, 4(29):48–50, January/April 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981976.

Anonymous:1844:SMMd

[Ano44r] Anonymous. Stated meeting, May 17. Proceedings of the American Philosophical Society, 4(30):82–85, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982203.

Anonymous:1844:SMMc

[Ano44s] Anonymous. Stated meeting, May 3. Proceedings of the American Philosophical Society, 4(30):67–82, April/June 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982202.

Anonymous:1844:SMNa

[Ano44t] Anonymous. Stated meeting, Nov. 1. Proceedings of the American Philosophical Society, 4(31):106–110, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982086.

Anonymous:1844:SMNb

[Ano44u] Anonymous. Stated meeting, November 15. Proceedings of the American Philosophical Society, 4(31):110–115, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982087.

Anonymous:1844:SMOb

[Ano44v] Anonymous. Stated meeting, October 18. Proceedings of the American Philosophical Society, 4(31):104–106, July/December 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982085.

Anonymous:1844:SMOa

[Ano44w] Anonymous. Stated meeting, October 4. Proceedings of the American Philosophical Society, 4(31):103–104, July/December 1844. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982084.

Anonymous:1844:SMS

[Ano44x] Anonymous. Stated meeting, September 20. Proceedings of the American Philosophical Society, 4(31):99–102, July/December 1844. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982083.

Anonymous:1845:Aa

[Ano45a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(32):??, January/April 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1845:Ab

[Ano45b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(33):??, May/August 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1845:Ac

[Ano45c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(34):??, September/December 1845. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1845:SMDb

[Ano45d] Anonymous. Special meeting, December 16. Proceedings of the American Philosophical Society, 4(34):219, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981841.

Anonymous:1845:SMFb

[Ano45e] Anonymous. Special meeting, February 14. Proceedings of the American Philosophical Society, 4(32):142, January/April 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981900.

Anonymous:1845:SMSa

[Ano45f] Anonymous. Special meeting, September 5. Proceedings of the American Philosophical Society, 4(34):189, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981834.

Anonymous:1845:SMAb

[Ano45g] Anonymous. Stated meeting, April 18. Proceedings of the American Philosophical Society, 4(32):153–156, January/April 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981905.

Anonymous:1845:SMAa

[Ano45h] Anonymous. Stated meeting, April 4. Proceedings of the American Philosophical Society, 4(32):151–153, January/April 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981904.

Anonymous:1845:SMAc

[Ano45i] Anonymous. Stated meeting, August 15. Proceedings of the American Philosophical Society, 4(33):184–188, May/August 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981847.

Anonymous:1845:SMDc

[Ano45j] Anonymous. Stated meeting, December 19. Proceedings of the American Philosophical Society, 4(34):219–223, September/ December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981842.

Anonymous:1845:SMDa

[Ano45k] Anonymous. Stated meeting, December 5. Proceedings of the American Philosophical Society, 4(34):214–219, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981840.

Anonymous:1845:SMFc

[Ano45l] Anonymous. Stated meeting, February 21. Proceedings of the American Philosophical Society, 4(32):143–146, January/April 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981901.

Anonymous:1845:SMFa

[Ano45m] Anonymous. Stated meeting, February 7. Proceedings of the American Philosophical Society, 4(32):138–142, January/April 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981899.

Anonymous:1845:SMJb

[Ano45n] Anonymous. Stated meeting, January 17. Proceedings of the American Philosophical Society, 4(32):135–138, January/April 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981898.

Anonymous:1845:SMJa

[Ano45o] Anonymous. Stated meeting, January 3. Proceedings of the American Philosophical Society, 4(32):133–135, January/April 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981897.

Anonymous:1845:SMJd

[Ano45p] Anonymous. Stated meeting, July 18. Proceedings of the American Philosophical Society, 4(33):181–183, May/August 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981846.

Anonymous:1845:SMJc

[Ano45q] Anonymous. Stated meeting, June 20. Proceedings of the American Philosophical Society, 4(33):170–181, May/August 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981845.

Anonymous:1845:SMMb

[Ano45r] Anonymous. Stated meeting, March 21. Proceedings of the American Philosophical Society, 4(32):149–151, January/April 1845. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981903.

Anonymous:1845:SMMa

[Ano45s] Anonymous. Stated meeting, March 7. Proceedings of the American Philosophical Society, 4(32):147–149, January/April 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981902.

Anonymous:1845:SMMd

[Ano45t] Anonymous. Stated meeting, May 16. Proceedings of the American Philosophical Society, 4(33):161–169, May/August 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981844.

Anonymous:1845:SMMc

[Ano45u] Anonymous. Stated meeting, May 2. Proceedings of the American Philosophical Society, 4(33):157–161, May/August 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981843.

Anonymous:1845:SMNa

[Ano45v] Anonymous. Stated meeting, Nov. 7. Proceedings of the American Philosophical Society, 4(34):203–210, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981838.

Anonymous:1845:SMNb

[Ano45w] Anonymous. Stated meeting, November 21. Proceedings of the American Philosophical Society, 4(34):210–214, September/ December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981839.

Anonymous:1845:SMOb

[Ano45x] Anonymous. Stated meeting, October 17. Proceedings of the American Philosophical Society, 4(34):198–202, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981837.

Anonymous:1845:SMOa

[Ano45y] Anonymous. Stated meeting, October 3. Proceedings of the American Philosophical Society, 4(34):197, September/December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981836.

Anonymous:1845:SMSb

[Ano45z] Anonymous. Stated meeting, September 19. Proceedings of the American Philosophical Society, 4(34):189–196, September/ December 1845. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981835.

Anonymous:1846:Aa

[Ano46a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(35):??, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1846:Ab

[Ano46b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(36):??, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1846:SMMe

[Ano46c] Anonymous. Special meeting, May 29. Proceedings of the American Philosophical Society, 4(35):254–255, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981383.

Anonymous:1846:SMAb

[Ano46d] Anonymous. Stated meeting, April 17. Proceedings of the American Philosophical Society, 4(35):248–250, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981380.

Anonymous:1846:SMAa

[Ano46e] Anonymous. Stated meeting, April 3. Proceedings of the American Philosophical Society, 4(35):245–248, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981379.

Anonymous:1846:SMAc

[Ano46f] Anonymous. Stated meeting, August 21. Proceedings of the American Philosophical Society, 4(36):275–278, July/December 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982091.

Anonymous:1846:SMDb

[Ano46g] Anonymous. Stated meeting, December 18. Proceedings of the American Philosophical Society, 4(36):296–298, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982098.

Anonymous:1846:SMDa

[Ano46h] Anonymous. Stated meeting, December 4. Proceedings of the American Philosophical Society, 4(36):295–296, July/December 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982097.

Anonymous:1846:SMFb

[Ano46i] Anonymous. Stated meeting, February 20. Proceedings of the American Philosophical Society, 4(35):237–240, January/June 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981376.

Anonymous:1846:SMFa

[Ano46j] Anonymous. Stated meeting, February 6. Proceedings of the American Philosophical Society, 4(35):233–237, January/June 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981375.

Anonymous:1846:SMJb

[Ano46k] Anonymous. Stated meeting, January 16. Proceedings of the American Philosophical Society, 4(35):227–232, January/June 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981374.

Anonymous:1846:SMJa

[Ano46l] Anonymous. Stated meeting, January 2. Proceedings of the American Philosophical Society, 4(35):225–227, January/June 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981373.

Anonymous:1846:SMJd

[Ano46m] Anonymous. Stated meeting, July 17. Proceedings of the American Philosophical Society, 4(36):271–275, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982090.

Anonymous:1846:SMJc

[Ano46n] Anonymous. Stated meeting, June 19. Proceedings of the American Philosophical Society, 4(35):255–269, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981384. See erratum [Ano47d].

Anonymous:1846:SMMb

[Ano46o] Anonymous. Stated meeting, March 20. Proceedings of the American Philosophical Society, 4(35):242–244, January/June 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981378.

Anonymous:1846:SMMa

[Ano46p] Anonymous. Stated meeting, March 6. Proceedings of the American Philosophical Society, 4(35):240–242, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981377.

Anonymous:1846:SMMc

[Ano46q] Anonymous. Stated meeting, May 1. Proceedings of the American Philosophical Society, 4(35):250–253, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981381.

Anonymous:1846:SMMd

[Ano46r] Anonymous. Stated meeting, May 15. Proceedings of the American Philosophical Society, 4(35):253–254, January/June 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981382.

Anonymous:1846:SMNb

[Ano46s] Anonymous. Stated meeting, November 20. Proceedings of the American Philosophical Society, 4(36):291–295, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982096.

Anonymous:1846:SMNa

[Ano46t] Anonymous. Stated meeting, November 6. Proceedings of the American Philosophical Society, 4(36):285–290, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982095.

Anonymous:1846:SMOb

[Ano46u] Anonymous. Stated meeting, October 16. Proceedings of the American Philosophical Society, 4(36):284–285, July/December 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982094.

Anonymous:1846:SMOa

[Ano46v] Anonymous. Stated meeting, October 2. Proceedings of the American Philosophical Society, 4(36):281–284, July/December 1846. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982093.

Anonymous:1846:SMS

[Ano46w] Anonymous. Stated meeting, September 18. Proceedings of the American Philosophical Society, 4(36):278–281, July/December 1846. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982092.

Anonymous:1847:Aa

[Ano47a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(37):??, January/March 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1847:Ab

[Ano47b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(38):??, April/June 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1847:Ac

[Ano47c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 4(39):??, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1847:ESM

[Ano47d] Anonymous. Erratum: Stated Meeting, June 19. Proceedings of the American Philosophical Society, 4(39):379, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982011. See [Ano46n].

Anonymous:1847:SMAb

[Ano47e] Anonymous. Stated meeting, April 16. Proceedings of the American Philosophical Society, 4(38):324–330, April/June 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982206.

Anonymous:1847:SMAa

[Ano47f] Anonymous. Stated meeting, April 2. Proceedings of the American Philosophical Society, 4(38):323–324, April/June 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982205.

Anonymous:1847:SMAc

[Ano47g] Anonymous. Stated meeting, August 20. Proceedings of the American Philosophical Society, 4(39):345–351, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982002.

Anonymous:1847:SMDb

[Ano47h] Anonymous. Stated meeting, December 17. Proceedings of the American Philosophical Society, 4(39):376–379, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982010.

Anonymous:1847:SMDa

[Ano47i] Anonymous. Stated meeting, December 3. Proceedings of the American Philosophical Society, 4(39):364–371, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982008.

Anonymous:1847:SMFb

[Ano47j] Anonymous. Stated meeting, February 19. Proceedings of the American Philosophical Society, 4(37):307–308, January/March 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981831.

Anonymous:1847:SMFa

[Ano47k] Anonymous. Stated meeting, February 5. Proceedings of the American Philosophical Society, 4(37):303–307, January/March 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981830.

Anonymous:1847:SMJa

[Ano47l] Anonymous. Stated meeting, January 1. Proceedings of the American Philosophical Society, 4(37):299–301, January/March 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981828.

Anonymous:1847:SMJb

[Ano47m] Anonymous. Stated meeting, January 15. Proceedings of the American Philosophical Society, 4(37):301–302, January/March 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981829.

Anonymous:1847:SMJd

[Ano47n] Anonymous. Stated meeting, July 16. Proceedings of the American Philosophical Society, 4(39):341–345, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982001.

Anonymous:1847:SMJc

[Ano47o] Anonymous. Stated meeting, June 18. Proceedings of the American Philosophical Society, 4(38):337–340, April/June 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982209.

Anonymous:1847:SMMb

[Ano47p] Anonymous. Stated meeting, March 19. Proceedings of the American Philosophical Society, 4(37):320–321, January/March 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981833.

Anonymous:1847:SMMa

[Ano47q] Anonymous. Stated meeting, March 5. Proceedings of the American Philosophical Society, 4(37):309–320, January/March 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981832.

Anonymous:1847:SMMd

[Ano47r] Anonymous. Stated meeting, May 21. Proceedings of the American Philosophical Society, 4(38):335–336, April/June 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982208.

Anonymous:1847:SMMc

[Ano47s] Anonymous. Stated meeting, May 7. Proceedings of the American Philosophical Society, 4(38):330–335, April/June 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982207.

Anonymous:1847:SMNb

[Ano47t] Anonymous. Stated meeting, November 19. Proceedings of the American Philosophical Society, 4(39):362–363, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982007.

Anonymous:1847:SMNa

[Ano47u] Anonymous. Stated meeting, November 5. Proceedings of the American Philosophical Society, 4(39):360–362, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982006.

Anonymous:1847:SMOa

[Ano47v] Anonymous. Stated meeting, October 1. Proceedings of the American Philosophical Society, 4(39):356–357, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982004.

Anonymous:1847:SMOb

[Ano47w] Anonymous. Stated meeting, October 15. Proceedings of the American Philosophical Society, 4(39):357–360, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982005.

Anonymous:1847:SMS

[Ano47x] Anonymous. Stated meeting, September 17. Proceedings of the American Philosophical Society, 4(39):351–356, July/December 1847. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982003.

Anonymous:1847:VI

[Ano47y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 4(39):381–388, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982000.

Anonymous:1848:Aa

[Ano48a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(40):??, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1848:Ab

[Ano48b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(41):??, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1848:FM

[Ano48c] Anonymous. Front matter. Proceedings of the American Philosophical Society, 5(40):??, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143363.

Anonymous:1848:SMOb

[Ano48d] Anonymous. Special meeting, October 13. Proceedings of the American Philosophical Society, 5(41):39, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143539.

Anonymous:1848:SMAb

[Ano48e] Anonymous. Stated meeting, April 21. Proceedings of the American Philosophical Society, 5(40):17–22, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143371.

Anonymous:1848:SMAa

[Ano48f] Anonymous. Stated meeting, April 7. Proceedings of the American Philosophical Society, 5(40):13–17, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143370.

Anonymous:1848:SMAc

[Ano48g] Anonymous. Stated meeting, August 18. Proceedings of the American Philosophical Society, 5(41):30–33, May/December 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143536.

Anonymous:1848:SMDa

[Ano48h] Anonymous. Stated meeting, December 1. Proceedings of the American Philosophical Society, 5(41):44–46, May/December 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143543.

Anonymous:1848:SMDb

[Ano48i] Anonymous. Stated meeting, December 15. Proceedings of the American Philosophical Society, 5(41):46–48, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143544.

Anonymous:1848:SMFb

[Ano48j] Anonymous. Stated meeting, February 18. Proceedings of the American Philosophical Society, 5(40):9–10, January/April 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143367.

Anonymous:1848:SMFa

[Ano48k] Anonymous. Stated meeting, February 4. Proceedings of the American Philosophical Society, 5(40):6–9, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143366.

Anonymous:1848:SMJb

[Ano48l] Anonymous. Stated meeting, January 21. Proceedings of the American Philosophical Society, 5(40):4–6, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143365.

Anonymous:1848:SMJa

[Ano48m] Anonymous. Stated meeting, January 7. Proceedings of the American Philosophical Society, 5(40):1–4, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143364.

Anonymous:1848:SMJd

[Ano48n] Anonymous. Stated meeting, July 21. Proceedings of the American Philosophical Society, 5(41):28–30, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143535.

Anonymous:1848:SMJc

[Ano48o] Anonymous. Stated meeting, June 16. Proceedings of the American Philosophical Society, 5(41):26–28, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143534.

Anonymous:1848:SMMb

[Ano48p] Anonymous. Stated meeting, March 17. Proceedings of the American Philosophical Society, 5(40):12–13, January/April 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143369.

Anonymous:1848:SMMa

[Ano48q] Anonymous. Stated meeting, March 3. Proceedings of the American Philosophical Society, 5(40):10–11, January/April 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143368.

Anonymous:1848:SMMd

[Ano48r] Anonymous. Stated meeting, May 19. Proceedings of the American Philosophical Society, 5(41):24–25, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143533.

Anonymous:1848:SMMc

[Ano48s] Anonymous. Stated meeting, May 5. Proceedings of the American Philosophical Society, 5(41):23–24, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143532.

Anonymous:1848:SMNb

[Ano48t] Anonymous. Stated meeting, November 17. Proceedings of the American Philosophical Society, 5(41):43–44, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143542.

Anonymous:1848:SMNa

[Ano48u] Anonymous. Stated meeting, November 3. Proceedings of the American Philosophical Society, 5(41):42–43, May/December 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143541.

Anonymous:1848:SMOc

[Ano48v] Anonymous. Stated meeting, October 20. Proceedings of the American Philosophical Society, 5(41):39–42, May/December 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143540.

Anonymous:1848:SMOa

[Ano48w] Anonymous. Stated meeting, October 6. Proceedings of the American Philosophical Society, 5(41):37–39, May/December 1848. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143538.

Anonymous:1848:SMS

[Ano48x] Anonymous. Stated meeting, September 15. Proceedings of the American Philosophical Society, 5(41):33–37, May/December 1848. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143537.

Anonymous:1849:Aa

[Ano49a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(42):??, January/March 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1849:Ab

[Ano49b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(43):??, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1849:SMAb

[Ano49c] Anonymous. Stated meeting, April 20. Proceedings of the American Philosophical Society, 5(43):83–86, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143441.

Anonymous:1849:SMAa

[Ano49d] Anonymous. Stated meeting, April 6. Proceedings of the American Philosophical Society, 5(43):81–83, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143440.

Anonymous:1849:SMAc

[Ano49e] Anonymous. Stated meeting, August 17. Proceedings of the American Philosophical Society, 5(43):98–99, April/September 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143446.

Anonymous:1849:SMFb

[Ano49f] Anonymous. Stated meeting, February 16. Proceedings of the American Philosophical Society, 5(42):70–72, January/March 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143499.

Anonymous:1849:SMFa

[Ano49g] Anonymous. Stated meeting, February 2. Proceedings of the American Philosophical Society, 5(42):67–70, January/March 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143498.

Anonymous:1849:SMJb

[Ano49h] Anonymous. Stated meeting, January 19. Proceedings of the American Philosophical Society, 5(42):53–66, January/March 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143497.

Anonymous:1849:SMJa

[Ano49i] Anonymous. Stated meeting, January 5. Proceedings of the American Philosophical Society, 5(42):49–53, January/March 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143496.

Anonymous:1849:SMJd

[Ano49j] Anonymous. Stated meeting, July 20. Proceedings of the American Philosophical Society, 5(43):95–98, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143445.

Anonymous:1849:SMJc

[Ano49k] Anonymous. Stated meeting, June 15. Proceedings of the American Philosophical Society, 5(43):90–94, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143444.

Anonymous:1849:SMMb

[Ano49l] Anonymous. Stated meeting, March 16. Proceedings of the American Philosophical Society, 5(42):73–79, January/March 1849. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143501.

Anonymous:1849:SMMa

[Ano49m] Anonymous. Stated meeting, March 2. Proceedings of the American Philosophical Society, 5(42):72–73, January/March 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143500.

Anonymous:1849:SMMd

[Ano49n] Anonymous. Stated meeting, May 18. Proceedings of the American Philosophical Society, 5(43):88–90, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143443.

Anonymous:1849:SMMc

[Ano49o] Anonymous. Stated meeting, May 4. Proceedings of the American Philosophical Society, 5(43):87–88, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143442.

Anonymous:1849:SMS

[Ano49p] Anonymous. Stated meeting, September 21. Proceedings of the American Philosophical Society, 5(43):100–103, April/September 1849. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143447.

Haldeman:1847:CAH

[Hal47] S. S. Haldeman. Corrections and additions to his paper on the Longicornia of the United States. Proceedings of the American Philosophical Society, 4(39):371–376, July/December 1847. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982009.